# On the Degrees of Freedom of some Lasso procedures

### v1

@@ 1 Introduction
Regularization has become a cornerstone of modern statistical modeling, particularly in regression settings where estimators are defined by minimizing an objective function comprising a loss term and a penalty on coefficient norms. Among such methods, the Lasso estimator (tibshirani1996), which employs the $\ell_{1}$ norm, achieves variable selection by shrinking a subset of coefficients exactly to zero, thereby inducing sparsity. Extensions such as the Group Lasso (yuan:2006:glasso) generalize this framework to structured settings, applying $\ell_{2}$ penalties to predefined groups of coefficients to enable group-wise selection. While these methods excel at regularization, their uniform shrinkage mechanisms lead to over shrink also the signal, thus producing biased estimates of the non zero coefficients. The Adaptive Lasso (zou.2006) and its extensions address this limitation by incorporating data-driven weights derived from preliminary least squares estimates. Unlike its non-adaptive counterpart, which imposes uniform shrinkage, the Adaptive Lasso adapts penalization to individual coefficients, reducing bias and attaining oracle properties under suitable conditions. Similarly, the Adaptive Group Lasso (wangleng2008csda), combines group-structured regularization with coefficient-specific adaptation. A central concept in statistical modeling is the notion of degrees of freedom. For example, in a normal linear regression framework, let $\mathbf{X}$ be a matrix of covariates with $p$ columns, $\mathbf{y}$ be the response vector of size $n$, and $\bm{\beta}$ the $p$-dimensional vector of unknown regression coefficients with $\mathbf{y}=\mathbf{X}\bm{\beta}+\bm{\epsilon}$ and $\bm{\epsilon}\sim\text{N}(\bm{0},\sigma^{2}\mathbf{I}_{n})$. In these settings, the degrees of freedom can be thought as the number of “independent pieces of information” needed to estimate the $p$ unknown parameters. For simple least-squares estimates of $\bm{\beta}$ the degrees of freedom equal the number of variables $p$ implying that each estimated quantity “consumes” one degree of freedom. A general notion of degrees of freedom is given by efron_dof as follows
@@
where $\widehat{\mathbf{y}}=(\widehat{y}_{1},\ldots,\widehat{y}_{n})$ is the vector of predictions obtained from the model under consideration and $\sigma^{2}$ is the (unknown) variance of the response. Notably, both $\mathbf{y}$ and $\widehat{\mathbf{y}}$ need to be considered as random quantities. In practice Equation (1) says that the more the predictions of the model are near to observed values, the higher the degrees of freedom. This notion of similarity can be written also in term of divergence, leading to the celebrated Stein’s Lemma (stein1981sure):
@@
Under these settings, zou2007 proved that an unbiased estimate of the degrees of freedom of Lasso regression is the number of nonzero coefficients estimated by the model, often referred as the size of the active set. This can be surprising, since differently from (random) subset regression, Lasso selects nonzero coefficients using the response variable so one would expect a higher value for the degrees of freedom. Nevertheless, this additional use of data is perfectly balanced with the amount of shrinkage that Lasso applies to each coefficient. Note that this would not happen in subset selection, where the “best” subset of variables to use in a linear model is chosen accordingly to some criteria (see tibshirani2015ss). Similar results were found for the generalized Lasso penalty (tibshirani_taylor_2011; Chen_etal.2020jasa) and for the more challenging case $p>n$ (dossal2013dof_lasso_general; tibshirani_taylor.2012). This perfect trade-off between shrinkage and the use of the response variable does not arise for the Group Lasso estimator, as discussed by vaiter_etal.2017. In this work, the authors derive formulae for computing the degrees of freedom of the Group Lasso under both orthogonal and non-orthogonal designs, generally yielding values lower than the size of the active set (see also vaiter2012dof_grouplasso). The concept of degrees of freedom underpins parametric inference and model selection: information criteria such as the Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) are defined by combining a model fit measure and the degrees of freedom (safken_etal.2025bmtrk). Despite their advantages, adaptive methods lack a closed-form expression for their effective degrees of freedom. While such estimators exist for non-adaptive methods like the Lasso (zou2007), adaptive variants resist analogous characterization due to their data-dependent weights, which introduce nonlinear dependencies between coefficients and penalties. Furthermore, the simplicity of the degrees of freedom formula for the Lasso estimator has inadvertently encouraged the misguided practice of computing information criteria for arbitrary penalized methods by substituting the size of the active set for the true degrees of freedom. This flawed approach risks overly optimistic model selection, often leading to overfitting and biased parameter estimates. In this paper, we solve this issue by deriving the first unbiased estimator of the effective degrees of freedom for the Adaptive Lasso within Stein’s unbiased risk estimation framework (stein1981sure). Our estimator explicitly accounts for the adaptive penalty’s dependence on data-driven weights, regularization parameters, and coefficient signs. The proof of our main results follows similar arguments to those in zou2007, which examined the first-order conditions of the Lasso optimization problem. However, in Adaptive Lasso and related models, an additional term appears that does not vanish. The explicit derivation of this term’s gradient with respect to the data is obtained using matrix differentiation rules. Additionally, we re-derive the expression for the degrees of freedom of the Group Lasso, providing additional insights and characterizations by means of an application of the implicit function theorem, thus complementing yuan:2006:glasso and vaiter_etal.2017. Leveraging these results, we are also able to provide analogous derivations for the more challenging Adaptive Group Lasso estimator. By bridging this theoretical gap, our work enables principled model selection, risk estimation, and complexity-aware inference for adaptive penalized regression methods.
@@ 2.1 The degrees of freedom of Adaptive Lasso
Let $\mathbf{X}$, $\mathbf{y}$, $\bm{\beta}$ and $\gamma$ as in the previous section, and $\mathbf{w}\in\mathbb{R}_{+}^{p}$ a vector of non-negative weights. Adaptive Lasso (zou.2006) estimator is defined as
@@
The set $\mathcal{A}$ is clearly dependent on the data and the values $\gamma$ but we omit this details for ease of notation. Calculation of the degrees of freedom involves the computation of the trace of $\partial\widehat{\mathbf{y}}_{\gamma}/\partial\mathbf{y}$. While for Lasso this result simplifies to the size of the active set (zou2007), for its adaptive version the presence of data-dependent weights makes its calculation less straightforward. We are now ready to present the first contribution of our paper, which is an unbiased estimator of the degrees of freedom for Adaptive Lasso model in Equation (4) for general expression of the adaptive weights.
@@
The conditions of Corollary 3 are typically satisfied in practice, as standard weights are usually monotonically decreasing functions of the absolute value of least squares estimates. The remaining two conditions, pertaining to the signs of the estimates and the size of the active set, are also frequently met in general settings and are guaranteed to hold for orthogonal designs. A counterintuitive property emerges from Corollary 3: within adjacent intervals defined by the transition points the estimated degrees of freedom $\widehat{df}_{\gamma}$ may be higher for $\gamma\in(\gamma_{l},\gamma_{l+1})$ than for $\gamma\in(\gamma_{l-1},\gamma_{l})$ even if the active set size is actually decreased. This implies that a larger $\gamma$ (and thus stronger regularization) does not always correspond to a simpler model, at least in terms of effective complexity, in line with the results of kaufman_rosset.2014 and janson_etal.2015. This apparent inconsistency arises because, as already discussed, $\widehat{df}_{\gamma}$ depends not only on the number of active parameters but also on their adaptive weighting structure. When $\gamma$ crosses a transition point $\gamma_{l}$, the adaptive weights—which themselves depend on $\mathbf{y}$—induce nonlinear interactions between the retained coefficients and the regularization path. These interactions can amplify the sensitivity of the fitted model to the data, effectively “consuming more information” despite a sparser active set.
@@ 2.2 The degrees of freedom of Group Lasso
Let $\mathbf{X}$, $\mathbf{y}$, $\bm{\beta}$ and $\gamma$ as in the previous section. Let $G\leq p\in\mathbb{N}$, be the number of groups and $\mathbf{w}\in\mathbb{R}^{G}_{+}$ a vector of non-negative weights associated to each group. The Group Lasso (yuan:2006:glasso) estimator is defined as
@@
For this problem the group weights $w_{g}$ are assumed to be independent from the response, a common choice being setting $w_{g}\propto n_{g}$ where $n_{g}$ is the size of the corresponding group. If the weights are determined in an adaptive way, then the model is known as Adaptive Group Lasso (wangleng2008csda) for which we provide results in the following Section 2.3. For the Group Lasso, let $\mathcal{G}_{p}=\{1,\dots,p\}$ be the index set of all variables and $\mathcal{G}_{G}=\{1,\dots,G\}$ the index set of all groups. Then, define the index set of active variables and the index set of active groups as
@@
The results in Equations (10) and (12) are apparently different from those of the Lasso and Adaptive Lasso. Furthermore, Equation (10) looks different from that reported in the original Group Lasso paper by yuan:2006:glasso and from a related expression reported by vaiter_etal.2017. The relations between these expressions are further discussed in Corollary S.3 in the Supplementary Materials. From Equations (10) and (12), it becomes evident that an unbiased estimate of the degrees of freedom for Group Lasso is again not merely the trace of the projection matrix—which would equate to the number of active variables—but rather the trace of a slightly different matrix. This matrix additionally depends on the design matrix and the vector of estimated coefficients. Consistently, using the size of the active set as an estimator of the degrees of freedom is generally not valid for Group Lasso. The following corollary provides bounds for the degrees of freedom obtained in Theorems 2 and 3.
@@ 2.3 The degrees of freedom of Adaptive Group Lasso
In the orthonormal design setting, an alternative expression for the estimated degrees of freedom appearing in Equation (15) in the spirit of yuan:2006:glasso can be derived, and it is presented in Corollary S.6 in the Supplementary Materials. In the same document, Corollary S.7 shows that Lasso and Group Lasso are special cases of Adaptive Group Lasso, and their estimated degrees of freedom can be recovered from Theorem 6.
@@ 3.2 Diabetes data
To illustrate the utility of our findings we analyze the Diabetes dataset (efron2004least; zou2007). This dataset is available in two versions: a small version with $p=10$ covariates, and a large version with $p=64$ covariates. The number of observations is $n=442$. The covariates are individual characteristics and biomarkers while the response variable is a quantitative measure of disease progression. We use this dataset to illustrate our results with two separate analyses. First, we use the original $p=10$ covariates of the small version and fit a linear model with Lasso and Adaptive Lasso penalties discussing model selection via BIC. In a second analysis we discretize the continuous covariates into ordinal categorical variables for “high”, “medium-high”, “medium-low”, and “low” levels of each biomarkers and encode these categorical variable into three dummy variables for each biomarker. After defining groups of dummy variable associated to the same categorical variable, we fit linear regression models with Group Lasso and Adaptive Group Lasso penalties. We start by discussing the results for the Lasso and Adaptive Lasso. The upper panels of Figure 2 report the estimated degrees of freedom as a function of $\log(\gamma)$. For comparison, we include the size of the active set (shown as a dashed line) to highlight discrepancies arising from the common but erroneous practice of equating this quantity with the estimated degrees of freedom across all penalization methods. It can be seen that for the Adaptive lasso, the correct estimated degrees of freedom (solid curve) are strictly lower-bounded by the active set size. This reveals that using the active set size as a proxy for degrees of freedom systematically underestimates the true complexity of the model, leading to
@@ 4 Conclusions
We introduced a general framework for computing unbiased estimates of the degrees of freedom in penalized regression models, extending the foundational work of zou2007. Our theoretical results underline that the common practice of using the size of active set as an estimate of the degrees of freedom is biased in many respect. This practice can severely distort inference since the size of active set and the true degrees of freedom are usually different, as highlighted by janson_etal.2015. For the Adaptive Lasso, for example, we demonstrated that the correct estimate includes, in addition to the active set size, an adjustment term whose sign depends on the weights’ selection. Under default weighting schemes, this term is positive, leading to an inflation of the degrees of freedom relative to the active set size. In contrast, the Group Lasso exhibits a deflation effect due to its reliance on the $\ell_{2}$ norm, reducing the degrees of freedom compared to the active set size. The Adaptive Group Lasso presents a more complex interplay: both inflation (from adaptive weights) and contraction (from $\ell_{2}$ shrinkage) coexist, making general characterization difficult. In addition to the analytical expressions for the degrees of freedom, we provide deep understanding of their local behavior as functions of $\gamma$. Specifically, we show that for the Adaptive Lasso (under certain conditions), these exhibit strictly positive but monotonic decreasing slopes within adjacent change points. For the Group Lasso, we show that the associated degrees of freedom are bounded between the active groups and active coefficients, and their local monotonicity is intimately related to the notion of orthogonality. These findings advance the understanding of model complexity in penalized regression, with implications for model selection, risk estimation, and theoretical analysis of high-dimensional methods.

### v2

@@ 1 Introduction
Regularization has become a cornerstone of modern statistical modeling, particularly in regression settings where estimators are defined by minimizing an objective function comprising a loss term and a penalty on coefficient norms. Among such methods, the Lasso estimator (Tibshirani, 1996), which employs the $\ell_{1}$ norm, achieves variable selection by shrinking a subset of coefficients exactly to zero, thereby inducing sparsity. Extensions such as the Group Lasso (Yuan and Lin, 2006) generalize this framework to structured settings, applying $\ell_{2}$ penalties to predefined groups of coefficients to enable group-wise selection. While these methods excel at regularization, their uniform shrinkage mechanisms lead to over shrink also the signal, thus producing biased estimates of the non zero coefficients. The Adaptive Lasso (Zou, 2006) and its extensions address this limitation by incorporating data-driven weights derived from preliminary least squares estimates. Unlike its non-adaptive counterpart, which imposes uniform shrinkage, the Adaptive Lasso adapts penalization to individual coefficients, reducing bias and attaining oracle properties under suitable conditions. Similarly, the Adaptive Group Lasso (Wang and Leng, 2008), combines group-structured regularization with coefficient-specific adaptation. A central concept in statistical modeling is the notion of degrees of freedom. For example, in a normal linear regression framework, let $\mathbf{X}$ be a matrix of covariates with $p$ columns, $\mathbf{y}$ be the response vector of size $n$, and $\bm{\beta}$ the $p$-dimensional vector of unknown regression coefficients with $\mathbf{y}=\mathbf{X}\bm{\beta}+\bm{\epsilon}$ and $\bm{\epsilon}\sim\text{N}(\bm{0},\sigma^{2}\mathbf{I}_{n})$. In these settings, the degrees of freedom can be thought as the number of “independent pieces of information” needed to estimate the $p$ unknown parameters. For simple least-squares estimates of $\bm{\beta}$ the degrees of freedom equal the number of variables $p$ implying that each estimated quantity “consumes” one degree of freedom. A general notion of degrees of freedom is given by Efron (1986) as follows
@@
where $\widehat{\mathbf{y}}=(\widehat{y}_{1},\ldots,\widehat{y}_{n})$ is the vector of predictions obtained from the model under consideration and $\sigma^{2}$ is the (unknown) variance of the response. Notably, both $\mathbf{y}$ and $\widehat{\mathbf{y}}$ need to be considered as random quantities. In practice Equation (1) says that the more the predictions of the model are near to observed values, the higher the degrees of freedom. This notion of similarity can be written also in term of divergence, leading to the celebrated Stein’s Lemma (Stein, 1981):
@@
Under these settings, Zou et al. (2007) proved that an unbiased estimate of the degrees of freedom of Lasso regression is the number of nonzero coefficients estimated by the model, often referred as the size of the active set. This can be surprising, since differently from (random) subset regression, Lasso selects nonzero coefficients using the response variable so one would expect a higher value for the degrees of freedom. Nevertheless, this additional use of data is perfectly balanced with the amount of shrinkage that Lasso applies to each coefficient. Note that this would not happen in subset selection, where the “best” subset of variables to use in a linear model is chosen accordingly to some criteria (see Tibshirani, 2015). Similar results were found for the generalized Lasso penalty (Tibshirani and Taylor, 2011; Chen et al., 2020) and for the more challenging case $p>n$ (Dossal et al., 2013; Tibshirani and Taylor, 2012). This perfect trade-off between shrinkage and the use of the response variable does not arise for the Group Lasso estimator, as discussed by Vaiter et al. (2017). In this work, the authors derive formulae for computing the degrees of freedom of the Group Lasso under both orthogonal and non-orthogonal designs, generally yielding values lower than the size of the active set (see also Vaiter et al., 2012). The concept of degrees of freedom underpins parametric inference and model selection: information criteria such as the Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) are defined by combining a model fit measure and the degrees of freedom (Säfken et al., 2024). Despite their advantages, adaptive methods lack a closed-form expression for their effective degrees of freedom. While such estimators exist for non-adaptive methods like the Lasso (Zou et al., 2007), adaptive variants resist analogous characterization due to their data-dependent weights, which introduce nonlinear dependencies between coefficients and penalties. Furthermore, the simplicity of the degrees of freedom formula for the Lasso estimator has inadvertently encouraged the misguided practice of computing information criteria for arbitrary penalized methods by substituting the size of the active set for the true degrees of freedom. This flawed approach risks overly optimistic model selection, often leading to overfitting and biased parameter estimates. In this paper, we solve this issue by deriving the first unbiased estimator of the effective degrees of freedom for the Adaptive Lasso within Stein’s unbiased risk estimation framework (Stein, 1981). Our estimator explicitly accounts for the adaptive penalty’s dependence on data-driven weights, regularization parameters, and coefficient signs. The proof of our main results follows similar arguments to those in Zou et al. (2007), which examined the first-order conditions of the Lasso optimization problem. However, in Adaptive Lasso and related models, an additional term appears that does not vanish. The explicit derivation of this term’s gradient with respect to the data is obtained using matrix differentiation rules. Additionally, we re-derive the expression for the degrees of freedom of the Group Lasso, providing additional insights and characterizations by means of an application of the implicit function theorem, thus complementing Yuan and Lin (2006) and Vaiter et al. (2017). Leveraging these results, we are also able to provide analogous derivations for the more challenging Adaptive Group Lasso estimator. By bridging this theoretical gap, our work enables principled model selection, risk estimation, and complexity-aware inference for adaptive penalized regression methods.
@@ 2.1 The degrees of freedom of Adaptive Lasso
Let $\mathbf{X}$, $\mathbf{y}$, $\bm{\beta}$ and $\gamma$ as in the previous section, and $\mathbf{w}\in\mathbb{R}_{+}^{p}$ a vector of non-negative weights. Adaptive Lasso (Zou, 2006) estimator is defined as
@@
The set $\mathcal{A}$ is clearly dependent on the data and the values $\gamma$ but we omit this details for ease of notation. Calculation of the degrees of freedom involves the computation of the trace of $\partial\widehat{\mathbf{y}}_{\gamma}/\partial\mathbf{y}$. While for Lasso this result simplifies to the size of the active set (Zou et al., 2007), for its adaptive version the presence of data-dependent weights makes its calculation less straightforward. We are now ready to present the first contribution of our paper, which is an unbiased estimator of the degrees of freedom for Adaptive Lasso model in Equation (4) for general expression of the adaptive weights.
@@
The conditions of Corollary 3 are typically satisfied in practice, as standard weights are usually monotonically decreasing functions of the absolute value of least squares estimates. The remaining two conditions, pertaining to the signs of the estimates and the size of the active set, are also frequently met in general settings and are guaranteed to hold for orthogonal designs. A counterintuitive property emerges from Corollary 3: within adjacent intervals defined by the transition points the estimated degrees of freedom $\widehat{df}_{\gamma}$ may be higher for $\gamma\in(\gamma_{l},\gamma_{l+1})$ than for $\gamma\in(\gamma_{l-1},\gamma_{l})$ even if the active set size is actually decreased. This implies that a larger $\gamma$ (and thus stronger regularization) does not always correspond to a simpler model, at least in terms of effective complexity, in line with the results of Kaufman and Rosset (2014) and Janson et al. (2015). This apparent inconsistency arises because, as already discussed, $\widehat{df}_{\gamma}$ depends not only on the number of active parameters but also on their adaptive weighting structure. When $\gamma$ crosses a transition point $\gamma_{l}$, the adaptive weights—which themselves depend on $\mathbf{y}$—induce nonlinear interactions between the retained coefficients and the regularization path. These interactions can amplify the sensitivity of the fitted model to the data, effectively “consuming more information” despite a sparser active set.
@@ 2.2 The degrees of freedom of Group Lasso
Let $\mathbf{X}$, $\mathbf{y}$, $\bm{\beta}$ and $\gamma$ as in the previous section. Let $G\leq p\in\mathbb{N}$, be the number of groups and $\mathbf{w}\in\mathbb{R}^{G}_{+}$ a vector of non-negative weights associated to each group. The Group Lasso (Yuan and Lin, 2006) estimator is defined as
@@
For this problem the group weights $w_{g}$ are assumed to be independent from the response, a common choice being setting $w_{g}\propto n_{g}$ where $n_{g}$ is the size of the corresponding group. If the weights are determined in an adaptive way, then the model is known as Adaptive Group Lasso (Wang and Leng, 2008) for which we provide results in the following Section 2.3. For the Group Lasso, let $\mathcal{G}_{p}=\{1,\dots,p\}$ be the index set of all variables and $\mathcal{G}_{G}=\{1,\dots,G\}$ the index set of all groups. Then, define the index set of active variables and the index set of active groups as
@@
The results in Equations (10) and (12) are apparently different from those of the Lasso and Adaptive Lasso. Furthermore, Equation (10) looks different from that reported in the original Group Lasso paper by Yuan and Lin (2006) and from a related expression reported by Vaiter et al. (2017). The relations between these expressions are further discussed in Corollary S.3 in the Supplementary Materials. From Equations (10) and (12), it becomes evident that an unbiased estimate of the degrees of freedom for Group Lasso is again not merely the trace of the projection matrix—which would equate to the number of active variables—but rather the trace of a slightly different matrix. This matrix additionally depends on the design matrix and the vector of estimated coefficients. Consistently, using the size of the active set as an estimator of the degrees of freedom is generally not valid for Group Lasso. The following corollary provides bounds for the degrees of freedom obtained in Theorems 2 and 3.
@@ 2.3 The degrees of freedom of Adaptive Group Lasso
In the orthonormal design setting, an alternative expression for the estimated degrees of freedom appearing in Equation (15) in the spirit of Yuan and Lin (2006) can be derived, and it is presented in Corollary S.6 in the Supplementary Materials. In the same document, Corollary S.7 shows that Lasso and Group Lasso are special cases of Adaptive Group Lasso, and their estimated degrees of freedom can be recovered from Theorem 6.
@@ 3.2 Diabetes data
To illustrate the utility of our findings we analyze the Diabetes dataset (Efron et al., 2004; Zou et al., 2007). This dataset is available in two versions: a small version with $p=10$ covariates, and a large version with $p=64$ covariates. The number of observations is $n=442$. The covariates are individual characteristics and biomarkers while the response variable is a quantitative measure of disease progression. We use this dataset to illustrate our results with two separate analyses. First, we use the original $p=10$ covariates of the small version and fit a linear model with Lasso and Adaptive Lasso penalties discussing model selection via BIC. In a second analysis we discretize the continuous covariates into ordinal categorical variables for “high”, “medium-high”, “medium-low”, and “low” levels of each biomarkers and encode these categorical variable into three dummy variables for each biomarker. After defining groups of dummy variable associated to the same categorical variable, we fit linear regression models with Group Lasso and Adaptive Group Lasso penalties. We start by discussing the results for the Lasso and Adaptive Lasso. The upper panels of Figure 2 report the estimated degrees of freedom as a function of $\log(\gamma)$. For comparison, we include the size of the active set (shown as a dashed line) to highlight discrepancies arising from the common but erroneous practice of equating this quantity with the estimated degrees of freedom across all penalization methods. It can be seen that for the Adaptive lasso, the correct estimated degrees of freedom (solid curve) are strictly lower-bounded by the active set size. This reveals that using the active set size as a proxy for degrees of freedom systematically underestimates the true complexity of the model, leading to
@@ 4 Conclusions
We introduced a general framework for computing unbiased estimates of the degrees of freedom in penalized regression models, extending the foundational work of Zou et al. (2007). Our theoretical results underline that the common practice of using the size of active set as an estimate of the degrees of freedom is biased in many respect. This practice can severely distort inference since the size of active set and the true degrees of freedom are usually different, as highlighted by Janson et al. (2015). For the Adaptive Lasso, for example, we demonstrated that the correct estimate includes, in addition to the active set size, an adjustment term whose sign depends on the weights’ selection. Under default weighting schemes, this term is positive, leading to an inflation of the degrees of freedom relative to the active set size. In contrast, the Group Lasso exhibits a deflation effect due to its reliance on the $\ell_{2}$ norm, reducing the degrees of freedom compared to the active set size. The Adaptive Group Lasso presents a more complex interplay: both inflation (from adaptive weights) and contraction (from $\ell_{2}$ shrinkage) coexist, making general characterization difficult. In addition to the analytical expressions for the degrees of freedom, we provide deep understanding of their local behavior as functions of $\gamma$. Specifically, we show that for the Adaptive Lasso (under certain conditions), these exhibit strictly positive but monotonic decreasing slopes within adjacent change points. For the Group Lasso, we show that the associated degrees of freedom are bounded between the active groups and active coefficients, and their local monotonicity is intimately related to the notion of orthogonality. These findings advance the understanding of model complexity in penalized regression, with implications for model selection, risk estimation, and theoretical analysis of high-dimensional methods.
Acknowledgements
The authors acknowledge support from the European Union - Next Generation EU, Mission 4 Component 2 via the MUR-PRIN grants - CUP C53D23002580006 ID 2022SMNNKY and CUP E53D23010290001 ID 2022KBTEBN. Mauro Bernardi also acknowledges partial funding by the BERN BIRD2222 01 - BIRD 2022 grant from the University of Padua.